OpenStudy (anonymous):

how do I simplify (3+5i)^4?

OpenStudy (anonymous):

just expand , you could get it into mod arg form and use demoivres theorm but that wont give exact answer

OpenStudy (anonymous):

ie do (3+5i)^2 (3+5i)^2 and expand, or use binomial theorm/pascals triangle expansion coeffiecents

OpenStudy (anonymous):

(3+5i)^2 (3+5i)^2 = -16+30i

OpenStudy (anonymous):

I got that answer...does that mean I do it twice?

OpenStudy (radar):

My guess is to expand taking (3+5i)(3+5i)=-16+30i then take (-16+30i)(-16+30) = -644-960i

OpenStudy (anonymous):

okay...that is what I was starting to do.

OpenStudy (radar):

Did you come up with the same results? I must admit the outcome does not look any simpler!!